18 September 2010 2nd half

LBC restart

  • Doug & Geno plugged in the cable that was brought to Tucson and repaired by Dave Baxter. All four chips on LBC red are working now. The bias level on chip 2 is high, due to a change made earlier. It should be restored to its original value.

  • Tonight's plan will be to:
    • (see below for tip/tilt to give to SX and DX).
    • check and correct pointing:
      • high elevation - run ptautoadjust which will automatically correct IE and CA.
      • low elevation to check IA.
      • back to high elevation.
    • Take more superfoc sequences to obtain Red/Blue and confirm any tilt or lack of tilt in focal plane. Created RB_rV_medcoarse_superfoc OBs with step size 0.06 mm, because analysis of blue data showed some variation in best focus and the minimum could be better measured if the sampling would go out a little further from focus.
    • Take data for the collimation model. Note that last night, to achieve co-pointing and to avoid limits at low elevation (30 deg) we added the following tip/tilt values:
      • to SX: tip = -37", tilt = +23"
      • to DX: tilt = -43"

  • 22:30 SW group wants to complete LBC tests in Binocular mode. I copied lbc to lbc_save20100918, then lbc_20100916 to lbc. lbckill/lbcstart. Restarted LBC.

  • Checking pointing. Steve clearing Active Optics, giving Z11=+1000 to SX, and giving above tip/tilt values to SX and DX. SX. Without first collimating, this put the DX at a limit.
  • DX out of collimation - needs manual adjustment. We give Z7 = Z8 = +1500nm. After it is nearly collimated, we are out of X,Y DX limits and can restore the DX tilt needed to copoint.

  • Slewed to pointing star ACT 0308, close to zenith. Collimated, Cap-S to stop because close. Observe Act0308. 062217/062213 B/R images
    • IE/CA
   Current IE = -27.802
   Current CA = -10.715
Reading header keywords:
  Coordinates: 23:44:00.9  29:17:48.2 

 Rotator Angle = -70.0664 Parallactic angle = -70.1250
 ROTX, ROTY:  1080.00 2913.00
Getting Star coordinates:
  Click on the Pointing Star...
2
 Star at pixel coordinates: 1111.806, 4096.583

  R.A.(-ve) Dec.Offsets: 7.124544, 265.122592

CA and IE Offsets: -246.8 -97.1

  NEW CA, IE: -257.5 -124.9
   Corrected IE = -124.888152109
   Corrected CA = -257.52469791456

Send commands to system to update pointing model? (n or N if no, any other character if yes)
y
Sending commands to system to update pointing model.
   ptmodify IE -124.888152109

Usage: utilityName [side] [term] [termValue]
         where utilityName = ptlist | ptreset | ptmodify
         side = 0 | 1 where 0=Left and 1=Right
         only ptmodify uses the parameters term and termValue

Exit 1
   ptmodify CA -257.52469791456

Usage: utilityName [side] [term] [termValue]
         where utilityName = ptlist | ptreset | ptmodify
         side = 0 | 1 where 0=Left and 1=Right
         only ptmodify uses the parameters term and termValue

Exit 1

...Done!
  • This pointing change was not accepted. We give up on correcting pointing for this test.
  • Steve slewed in binocular mode to bright star, then we want to offset 1800" North (to get bright star out of field), but it did not work.
  • Guided 180sec image is lbcr - 071328. Guide corrections are typically small in Az and El is +/- 0.2".
  • changed weighting
  • xxxx
  • changed weighting again
  • 072335 lost guiding
  • back to 50/50 and taking binocular observations. 072847/072852 R/B
  • changed weighting to 99.9 on right side
  • took bino image
  • blue weighted
  • 073753 r and
  • blue only: 074346.
  • 99.9 on the right. 074752
  • another weighting - 075202.
  • Done with TCS testing.

  • turn off camera. cp lbc_20100421 lbc to restore working LBC code for AO5. Steve changing TCS version to AO5. Calling Vincenzo et al. about bias offset to Red chip 2.
  • Examined this file which Vincenzo edited, but have not yet changed it.
  • Had cleared ao after running down/up tcs.
  • Need to give Z7=Z8=+1500nm on DX again.
  • Focus/collimate near ACT 0295. Well, dofpia crashed on Blue (which was nearly collimated). See 084945.
Not enough points for estimating Z22
Rejected pupil 3 with diameter 3155.53
% Attempt to subscript RES with <INT      (       3)> is out of range.
% Execution halted at: LBCFPIA           292 /home/lbcobs/LBCFPIA/lbcfpia/src/lbcfpia.pro
%                      DOFPIA            444 /home/lbcobs/LBCFPIA/lbcfpia/src/dofpia.pro
  • Backout, give tip/tilt corrections because we had hit an SX limit, and restart dofpia, shift-S to exit.
  • ACT0295 at elev=81 deg: 090537/090534 B/R ----> CA and IE Offsets: -33.8 3.5, NEW CA, IE: -44.5 -24.3
  • run mirradjust to check copointing at this high elevation. To SX: tip now -25" and tilt now +25"
  • take pair again: 091730/091723 B/R looks well-pointed and well-co-pointed. Elev=83deg.
  • Slew to low elevation to check range WT10_104 - this can be our first collimation model point. Tambient=12.0 C.
    • B/R 092748/092749
    • 1sec exposures in focus: 093015/093011 and then WT10_104 to check IE/CA and IA. Actually entered "no" and not "n" or "N" to ptautoadjust so I inadvertantly changed the pointing.
  • Slew to another pointing star at moderate elevation. WT10_278. Focus/collimate.
  • Another collimation point: elevation=76.
    • B/R 094450/094451
    • 1sec exposure 094659/094655
  • Take 2 superfoc sequences here to double-check fp tilt of blue and red (medcoarse). 095928
    • PA=0: RB_rV_medcoarse
    • PA=180: RB_rV_medcoarse_superfoc
  • Check pointing here at WT10_278 ... after this done with pointing checks. (will measure rotator center & collect more collim data) 095918
    • CA and IE Offsets: 15.0 15.7; NEW CA, IE: -45.1 -37.2
  • Elevation 57 deg. 100954/100956 B/R and 1sec exposure 101153/101148
  • Elevation 47 deg: 102148/102149 B/R and 1sec exp 102345/102340
  • Elevation 37 deg: 1sec exp 103202/103158
  • Elevation 27 deg: 1sec exp 104340/104336
  • Elevation 64 deg: 105506/105505 1sec exp 105733/105728
  • Elevation 74 deg: 110224/110225 1sec exp 110505/110501
  • Elevation 78 deg: 111222/111223 1sec exp 111418/111413
  • Elevation 85.8 deg: Red converged early but Blue did not... after 8 iterations, Shift-S and restart: 113319/113318 (82.3deg) 113518/113513
  • Slew to 69 deg - spontaneously we had warnings for both blue/red and many readings dropped off the User Interface... instead just dashes. Was at or just before about 11:38 UT. But it all came back up by itself.
  • Elevation 69 deg: LBC restarted after a crash... at 11:48 UT. Filed issue #2844 about spontaneous LBC crashes. Resuming Blue is very slow to converge. 120523/120522 1sec exp 120721/120716.
  • PA=180 for rotator center check using previous 1sec images.
  • getting light - try superfoc at PA=180 again. And another at PA=0.
  • a few biases before resetting chip 2 offset, but these are being done with dome open and telescope slewing as we close up. 121854/121907... 122111/122118.
  • Turning off LBC. For Red: bias levels are approx 325/1444/298/310 with current offsets. Turned off LBC and changed Chip 2 offset (2.4805 to 2.5805).
    • 123838/123848/ now bias levels are: 349/151/320/343
    • Changed again, now to 2.5661. Restart LBC. and take series of biases (note lights are on in dome).
    • Bias levels now look well matched: 328/328/302/314

  • Summary We took some TCS time to plug in the cable and then ran a little longer than midnight in completing some TCS tests. We obtained data to check the rotator center, tilt, and to use to generate collimation lookup tables for both Blue and Red.

-- OlgaKuhn - 18 Sep 2010
Topic revision: r3 - 18 Sep 2010, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback